# Range Image-Based Implicit Neural Compression for LiDAR Point Clouds

### v1

@@ III-F RI Decoder
The depth image reconstruction is in a two-stage manner. Each patch is first reconstructed by feeding the sets of pairs of a patch index and an in-patch coordinate $\{(i,i_{u},i_{v})\ |\ i\in\{1,\cdots,N_{p}^{2}\},i_{u}\in\{1,\cdots,W\},i_{v}\in\{1,\cdots,H\}\}$ to the depth INR The reconstructed patches are then gathered to build the complete depth image $\hat{I}_{D}$. Finally, the reconstructed RI $\hat{I}$ is obtained as
$\hat{I}(u,v)=\begin{cases}\hat{I}_{D}(u,v)&\hat{I}_{M}(u,v)=0,\\ \rho_{\rm null}&\mathrm{otherwise}.\end{cases}$ (13)
@@
The concluding phase of our decoding procedure is the reconstruction of a 3D point cloud through a 2D-to-3D mapping against the reconstructed RI. When the pixel $(u,v)$ on the RI has a valid point depth, i.e., is not $\rho_{\rm null}$, the corresponding point in the spherical coordinate $\hat{\mathbf{p}}^{\prime}=(\hat{\rho},\hat{\phi},\hat{\theta})$ is obtained as follows:
@@
Finally, the 3D points in the spherical coordinate are transformed into the 3D Cartesian coordinate $\hat{\mathbf{p}}=(\hat{x},\hat{y},\hat{z})$ as
@@ IV-A Settings
Dataset: We use the KITTI dataset [10] as our source of 3D point cloud data. For R-D performance, we evaluate the KITTI Odometry dataset, using frames 00, 25, 50, 75, and 100 from sequences 00 to 06. For downstream tasks, we use the KITTI 3D Object Detection dataset, specifically evaluating frame 000134 to analyze the impact of compression on object detection accuracy. All data were collected with a Velodyne 64 scanner that features 64 laser scan lines and an azimuth resolution of 0.09 degrees.
@@
Figure 3: Chamfer distance as a function of bitrates across different sequences of KITTI’s LiDAR point cloud. From left to right: (left) Seq:00-00, (middle) Seq:00-25, and (right) Seq:00-50.
Refer to caption (a) Seq:00-00 Refer to caption (b) JPEG
bpp:2.89
CD:0.717
Refer to caption (c) JPEG2000
bpp:2.55
CD:0.436
Refer to caption (d) HEIF
bpp:2.35
CD:0.339
Refer to caption (e) AVIF
bpp:1.41
CD:0.324
Refer to caption (f) COIN
bpp:3.02
CD:0.991
Refer to caption (g) G-PCC
bpp:2.62
CD:0.137
Refer to caption (h) Draco
bpp:2.81
CD:0.127
Refer to caption (i) R-PCC (Deflate)
bpp:2.73
CD:0.0867
Refer to caption (j) R-PCC (LZ4)
bpp:3.63
CD:0.229
Refer to caption (k) Proposed
(pixel-wise)
bpp:2.99
CD:0.123
Refer to caption (l) Proposed
bpp:2.55
CD:0.0902
Refer to caption (m) Seq:00-25 Refer to caption (n) JPEG
bpp:2.47
CD:1.01
Refer to caption (o) JPEG2000
bpp:2.55
CD:0.441
Refer to caption (p) HEIF
bpp:2.67
CD:0.323
Refer to caption (q) AVIF
bpp:1.51
CD:0.308
Refer to caption (r) COIN
bpp:3.02
CD:0.957
Refer to caption (s) G-PCC
bpp:2.30
CD:0.136
Refer to caption (t) Draco
bpp:2.53
CD:0.128
Refer to caption (u) R-PCC (Deflate)
bpp:2.48
CD:0.106
Refer to caption (v) R-PCC (LZ4)
bpp:4.05
CD:0.235
Refer to caption (w) Proposed
(pixel-wise)
bpp:2.04
CD:0.107
Refer to caption (x) Proposed
bpp:2.55
CD:0.0982
Figure 4: Snapshots of the reconstructed LiDAR point clouds in proposed and baseline methods. Here, (b)-(l) and (n)-(x) show the reconstructed point clouds of sequences 00-00 and 00-25, respectively.
TABLE I: The list of BD-CD $\uparrow$ for KITTI dataset for the different sequences.
Seq. JPEG JPEG2000 HEIF AVIF COIN G-PCC Draco
R-PCC
R-PCC
00-00 0.887 0.426 0.266 0.149 0.955 0.054 0.054 0.012 0.124
00-25 0.924 0.435 0.263 0.137 0.894 0.029 0.040 0.007 0.146
00-50 1.198 0.531 0.294 0.166 1.050 0.024 0.025 0.014 0.156
00-75 1.261 0.563 0.309 0.169 0.905 0.022 0.044 -0.019 0.091
00-100 0.806 0.383 0.232 0.130 1.019 0.036 0.040 -0.026 0.096
01-00 0.825 0.391 0.242 0.109 0.990 0.041 0.064 -0.039 0.101
01-25 0.753 0.385 0.250 0.142 0.999 0.100 0.113 -0.017 0.114
01-50 0.822 0.393 0.259 0.147 1.013 0.065 0.086 0.000 0.120
01-75 1.072 0.476 0.287 0.153 0.897 0.053 0.060 -0.004 0.102
01-100 1.037 0.485 0.281 0.170 1.001 0.045 0.074 0.014 0.156
02-00 0.996 0.441 0.259 0.121 0.902 0.015 0.032 -0.044 0.137
02-25 0.769 0.380 0.236 0.129 0.919 0.036 0.074 -0.011 0.110
02-50 0.775 0.384 0.234 0.143 0.933 0.067 0.082 -0.024 0.094
02-75 0.743 0.369 0.249 0.149 0.882 0.036 0.095 -0.022 0.085
02-100 0.838 0.389 0.250 0.170 1.063 0.057 0.096 -0.008 0.095
03-00 1.065 0.440 0.251 0.139 0.927 0.021 0.021 0.012 0.162
03-25 1.241 0.592 0.319 0.118 0.873 -0.006 -0.012 -0.026 0.106
03-50 0.905 0.409 0.244 0.125 0.975 0.013 0.017 -0.052 0.092
03-75 0.816 0.364 0.219 0.109 0.856 0.028 0.034 -0.067 0.054
03-100 0.810 0.369 0.227 0.122 1.025 0.040 0.067 -0.054 0.077
04-00 0.769 0.352 0.215 0.127 0.987 0.031 0.036 -0.027 0.159
04-25 1.037 0.502 0.265 0.083 0.954 0.006 0.013 -0.042 0.105
04-50 1.103 0.476 0.272 0.147 1.008 0.024 0.047 -0.006 0.121
04-75 0.919 0.401 0.232 0.131 0.858 0.005 0.011 -0.038 0.104
04-100 0.954 0.435 0.263 0.162 1.079 0.047 0.075 -0.003 0.102
05-00 0.807 0.402 0.240 0.144 0.942 0.036 0.038 0.008 0.149
05-25 0.811 0.396 0.243 0.168 0.929 0.042 0.052 -0.024 0.081
05-50 1.271 0.576 0.300 0.131 0.920 0.001 0.018 -0.037 0.089
05-75 1.380 0.634 0.333 0.118 0.852 -0.013 -0.014 -0.043 0.106
05-100 1.064 0.467 0.279 0.170 1.007 0.017 0.039 -0.030 0.084
06-00 0.836 0.388 0.219 0.093 0.868 0.010 0.003 -0.034 0.146
06-25 1.085 0.477 0.246 0.073 0.864 -0.001 0.001 -0.058 0.094
06-50 1.015 0.459 0.238 0.076 0.871 -0.007 -0.020 -0.064 0.080
06-75 1.256 0.612 0.321 0.078 0.830 -0.013 -0.013 -0.046 0.099
06-100 0.793 0.374 0.213 0.075 0.963 0.025 0.034 -0.065 0.080
Average 0.961 0.444 0.259 0.131 0.943 0.028 0.041 -0.025 0.109
TABLE II: 3D object detection accuracy $\uparrow$ for different model sizes.
IoU G-PCC R-PCC Proposed
Type Object Low Mid High Low Mid High Low Mid High
2D IoU Pedestrian 0.07 0.27 0.68 0.08 0.16 0.18 0.43 0.45 0.52
Cyclist 0.22 0.70 0.83 0.30 0.55 0.58 0.86 0.90 0.82
Car - 0.19 0.20 0.30 0.26 0.22 0.22 0.23 0.26
BEV IoU Pedestrian - 0.04 0.31 - 0.04 0.02 0.74 0.67 0.57
Cyclist - 0.15 0.81 0.03 0.15 0.16 0.72 0.71 0.73
Car - - - - - - - - -
3D IoU Pedestrian - 0.04 0.28 - 0.04 0.02 0.54 0.51 0.50
Cyclist - 0.15 0.75 0.02 0.142 0.142 0.64 0.68 0.67
Car - - - - - - - - -
TABLE III: Average decoding latency $\downarrow$
Method Latency per sequence (ms)
JPEG 0.49
JPEG2000 0.54
HEIF 0.51
AVIF 0.48
COIN 0.71
G-PCC 2.00
Draco 3.00
R-PCC (Deflate) 11.5
R-PCC (LZ4) 10.3
Proposed 0.69
Metric: Regarding the metrics for the decoded 3D point clouds, we follow the common practice in the community using chamfer distance (CD). CD is defined as:
@@
For the R-D performance assessment between the proposed method and the baselines, we use the Bjøntegaard delta chamfer distance (BD-CD) [38] for calculating average CD improvement between R-D curves for the same bitrate, where positive values denote CD improvement compared to the baselines.
@@
The input to the depth INR is a concatenation of the patch index $i$ and the in-patch pixel coordinates $(i_{u},i_{v})\in\mathbb{R}^{2}$, resulting in a 3-dimensional input vector. Since we set the patch scaling factor to $N_{p}=16$, the patch index $i$ ranges from 00 to $255$. The architecture of the depth INR is as follows:
@@
    We also select Draco [39] as the 3D point cloud compression baseline which also belongs to the PCC family. We use the official implementation of the Draco encoder that performs KD-tree-based compression [40].
@@
  4. 4)
    R-PCC [22] is an RI based LiDAR compression baseline. It maps LiDAR point clouds to RI s and performs intra-coding using floating-point lossless coding methods. Here, we use LZ4 and Deflate for coding methods due to their fast decompression.
@@
    COmpression with Implicit Neural representations (COIN) [7] is an INR–based image compression baseline. The INR architecture is trained to obtain a direct mapping of the pixel coordinate to each pixel value of RI. We assume that COIN serves as a reliable indicator of the efficiency of our depth/mask separation strategy, as it does not employ the process.
@@ IV-B1 Rate-Distortion Performance
  • The proposed method achieves the best 3D reconstruction quality in the bpp range of 1 through 2.5.
@@
Figs. 4 (a)-(x) show the snapshots of the original and reconstructed LiDAR point clouds in each method at approximately 2.55 bpp regimes. Here, Figs. 4 (a)-(l) and (m)-(x) use the sequences 00-00 and 00-25, respectively. The proposed method can reconstruct a clean point cloud at the same bitrate. However, all PCC, image compression, and INR-based compression methods contain circular noises and/or decrease the number of 3D points in the reconstructed LiDAR point clouds. A circular noise still remains in R-PCC methods as well.
Table I lists the BD-CD of the proposed method against the baselines in each sequence of LiDAR point clouds. It shows that the proposed method achieves the best 3D reconstruction quality in the same bitrate range against PCC, image compression, and INR-based compression methods irrespective of LiDAR sequences. In addition, the performance gap between R-PCC and proposed methods depends on the lossless coding method. When R-PCC uses a low-efficiency coding method, such as LZ4, for fast decoding, the proposed scheme achieves better R-D performance than R-PCC.
@@ IV-B2 Downstream Task
We then discuss the impact of our RI compression method on the performance of downstream tasks on the LiDAR point cloud. We selected 3D object detection as a representative example of downstream tasks. We used PointPillar [41] as a 3D object detector, and we used our method and G-PCC and R-PCC (Deflate), which are baselines with better BD-CD performance, to compress the point clouds and fed the reconstructed ones into PointPillar. In addition, we assumed three model sizes (approximately bpp 1.5 as Low, 2.5 as Middle, and 3.5 as High) in each compression method. Table III shows the 2D Intersection over Union (IoU), Birds Eye View (BEV) IoU, and 3D IoU in each object for each compression approach. Especially in limited bpp scenarios, the proposed approach performs better than other baselines in most IoU metrics.
@@ IV-B3 Decoding Latency
Table III shows the average decoding latency of the proposed and baseline methods for LiDAR sequence of 00-00. The decoding latency values for the proposed method and the baselines are the total time required from RI decoding to the 2D-to-3D mapping. The decoding delay of the proposed method is comparable to that of image compression methods and has more than 65.5% and 93.3% reduction compared to the PCC and R-PCC methods, respectively. This means that the proposed method approaches the decoding delay of image compression methods and achieves 3D reconstruction quality comparable to PCC/R-PCC methods.
(a) R-D performance.
(b) Convergence speed.
@@
(a) CD for different pruning sparsity.
(b) Quantization with different $N_{b}$.
Figure 6: Model compression performance.

### v2

@@ III-F RI Decoder
The depth image reconstruction is in a two-stage manner. Each patch is first reconstructed by feeding the sets of pairs of a patch index and an in-patch coordinate $\{(i,i_{u},i_{v})\ |\ i\in\{1,\cdots,N_{p}^{2}\},i_{u}\in\{1,\cdots,W/N_{p}\},i_{v}\in\{1,\cdots,H/N_{p}\}\}$ to the depth INR. The reconstructed patches are then gathered to build the complete depth image $\hat{I}_{D}$. Finally, the reconstructed RI $\hat{I}$ is obtained as
$\hat{I}(u,v)=\begin{cases}\hat{I}_{D}(u,v)&\hat{I}_{M}(u,v)=1,\\ \rho_{\mathrm{null}}&\mathrm{otherwise}.\end{cases}$ (13)
@@
[…]
@@
[…]
@@ IV-A Settings
Dataset: We use the KITTI dataset [kitti] as our source of 3D point cloud data. For R-D performance, we evaluate the KITTI Odometry dataset, using frames 00, 25, 50, 75, and 100 from sequences 00 to 06. For downstream tasks, we use the KITTI 3D Object Detection dataset. We split the official training set into 3,712 training samples and 3,769 validation samples, and evaluate detection performance on the validation split to quantify the impact of compression. We use OpenPCDet [openpcdet2020] v0.6.0 to train and evaluate three representative detectors: PointPillars [lang2019pointpillars], SECOND [yan2018second], and PointRCNN [shi2019pointrcnn].
All data were collected with a Velodyne 64 scanner that features 64 laser scan lines and an azimuth resolution of 0.09 degrees. In the proposed scheme, these 3D point clouds are projected into an RI for compression. Note that projecting points into a range image may cause point loss. We evaluate the point retention ratio defined as $r_{N}=|\hat{\mathcal{P}}|/|\mathcal{P}|$, where $\mathcal{P}$ is the original input point cloud and $\hat{\mathcal{P}}$ is the reconstructed point cloud via RI projection and back-projection. On the KITTI odometry dataset, we observe $r_{N}=41.02\%\pm 0.38\%$ using 35 selected frames from sequences 00–06. Similarly, we observe $r_{N}=74.33\%\pm 0.69\%$ on the KITTI object detection validation split of 3,769 frames, measured on the cropped point clouds used as inputs for 3D object detectors.
@@
Figure 3: Chamfer distance as a function of bitrate across different sequences of the KITTI LiDAR point clouds, where the bitrate is measured in bits per point (bpp). From left to right: (left) performance in frame 00 of sequence 00, (middle) performance in frame 25 of sequence 00, and (right) performance in frame 50 of sequence 00.
Metric: Regarding the metrics for the decoded 3D point clouds, we follow the common practice in the community using chamfer distance (CD). CD has been widely adopted as a distortion measure for 3D point cloud reconstruction and sensing [REF01].
CD is defined as:
@@
For the R-D performance assessment between the proposed method and the baselines, we use the Bjøntegaard delta chamfer distance (BD-CD) [bjotegaard2001calculation] for calculating average CD improvement between R-D curves for the same bitrate, where positive values denote CD improvement compared to the baselines. For downstream tasks, we evaluate 3D object detection accuracy using the Car 3D bounding-box average precision (AP).
@@
The input to the depth INR is a concatenation of the patch index $i$ and the in-patch pixel coordinates $(i_{u},i_{v})\in\mathbb{R}^{2}$, resulting in a 3-dimensional input vector. Since we set the patch scaling factor to $N_{p}=16$, the patch index $i$ ranges from 0 to $255$. The architecture of the depth INR is as follows:
@@
    We also select Draco [bib:draco] as the 3D point cloud compression baseline which also belongs to the PCC family. We use the official implementation of the Draco encoder that performs KD-tree-based compression [kd-tree].
@@
    OctAttention [OctAttention] is an octree-based autoencoder within the PCC family. This method improves the conventional octree structure by incorporating attention mechanisms for better context modeling. To evaluate its performance across different compression levels, we set the octree depth to values from 8 to 13.
  4. 4)
@@
    R-PCC [wang2022r] is an RI based LiDAR compression baseline. It maps LiDAR point clouds to RI s and performs intra-coding using floating-point lossless coding methods. Here, we use LZ4 and Deflate for coding methods due to their fast decompression.
  6. 6)
    COmpression with Implicit Neural representations (COIN) [dupont2021coin] is an INR–based image compression baseline. The INR architecture is trained to obtain a direct mapping of the pixel coordinate to each pixel value of RI. We assume that COIN serves as a reliable indicator of the efficiency of our depth/mask separation strategy, as it does not employ the process.
@@
Refer to caption ((a)) Refer to caption ((b)) Refer to caption ((c)) Refer to caption ((d)) Refer to caption ((e)) Refer to caption ((f))
Refer to caption ((g)) Refer to caption ((h)) Refer to caption ((i)) Refer to caption ((j)) Refer to caption ((k)) Refer to caption ((l))
Refer to caption ((m)) Refer to caption ((n)) Refer to caption ((o)) Refer to caption ((p)) Refer to caption ((q)) Refer to caption ((r))
Refer to caption $($s)) Refer to caption ((t)) Refer to caption ((u)) Refer to caption ((v)) Refer to caption ((w)) Refer to caption ((x))
Figure 4: Snapshot of the reconstructed LiDAR point clouds in proposed and baseline methods. Here, (b)-(l) and (n)-(x) show the reconstructed point clouds of frames 00 and 25 of sequence 00, respectively.
TABLE I: The list of BD-CD $\uparrow$ for the KITTI dataset across the different sequences. Note that BD-CD is evaluated for each baseline using the proposed method as the reference. Positive values indicate that the proposed method achieves a lower chamfer distance than the corresponding baseline.
[…]
Oct
Attention
$\S$
R-PCC
$\P$
R-PCC
$\P$
00 1.393 0.801 0.535 0.256 1.103 0.081 0.157 -0.025 0.023 0.110
01 1.234 0.686 0.515 0.259 1.095 0.097 0.234 -0.017 0.019 0.097
02 1.120 0.612 0.447 0.254 1.039 0.083 0.226 -0.026 0.006 0.094
03 1.322 0.733 0.497 0.232 1.074 0.059 0.176 -0.045 -0.007 0.078
04 1.343 0.748 0.515 0.237 1.092 0.074 0.192 -0.039 0.010 0.097
05 1.484 0.877 0.576 0.248 1.041 0.065 0.153 -0.040 0.008 0.095
06 1.324 0.785 0.527 0.224 0.993 0.053 0.153 -0.070 -0.024 0.062
Average 1.317 0.749 0.516 0.244 1.063 0.073 0.185 -0.037 0.005 0.090
  • $\dagger$: Image based method(s),  $\ddagger$}: INR based method,  $\S$: Point Cloud based method(s),  $\P$: Range Image based method(s).
((a))
((b))
((c))
Figure 5: Quantitative results of 3D object detection on the KITTI detection dataset. Car 3D bounding-box AP@0.7, 0.7, 0.7 as a function of bitrate (bits per point), for three downstream detectors (left to right): PointPillars, SECOND, and PointRCNN.
TABLE II: Average decoding latency $\downarrow$
Method Latency per frame
JPEG 0.49 ms
JPEG2000 0.54 ms
HEIF 0.51 ms
AVIF 0.48 ms
COIN 0.71 ms
G-PCC 2.00 ms
Draco 3.00 ms
OctAttention 10.6 s
R-PCC (Deflate) 11.5 ms
R-PCC (LZ4) 10.3 ms
Proposed 0.69 ms
TABLE III: Average encoding latency $\downarrow$
Method Latency per frame
JPEG 9 ms
JPEG2000 10 ms
HEIF 55 ms
AVIF 94 ms
COIN 30 min
G-PCC 65 ms
Draco 10 ms
OctAttention 134 ms
R-PCC (Deflate) 20 ms
R-PCC (LZ4) 60 ms
Proposed 180 min
@@ IV-B1 RATE-DISTORTION PERFORMANCE
  • The proposed method achieves higher 3D reconstruction quality than G-PCC, Draco, image compression, and RI compression methods across the bpp range up to 3.0 for frame 00.
  • OctAttention achieves the best R-D performance in frames 00, 25 and 50, whereas it requires long decoding latency, as will be detailed in Table II.
@@
Figs. 4 (a)-(x) show the snapshots of the original and reconstructed LiDAR point clouds in each method Here, Figs. 4 (a)-(l) and (m)-(x) use frames 00 and 25 of sequence 00, respectively. The proposed method can reconstruct a clean point cloud at the same bitrate. However, some PCC, image compression, and INR-based compression methods contain circular noises and/or decrease the number of 3D points in the reconstructed LiDAR point clouds. A circular noise still remains in R-PCC methods as well.
Table I lists the average BD-CD performance of the proposed method against the baselines in each sequence of LiDAR point clouds. Here, BD-CD is evaluated for each baseline using the proposed method as the reference. It shows that the proposed method achieves the best 3D reconstruction quality in the same bitrate range against G-PCC, Draco, image compression, and INR-based compression methods irrespective of LiDAR sequences. For OctAttention and R-PCC (Deflate), the proposed method can be comparable or slightly worse in some cases. In addition, the performance gap between R-PCC and the proposed methods depends on the lossless coding method. When R-PCC uses a low-efficiency coding method, such as LZ4, for fast decoding, the proposed scheme achieves better R-D performance than R-PCC.
@@ IV-B2 DOWNSTREAM TASK
We then discuss the impact of our RI compression method on the performance of downstream tasks on the LiDAR point cloud. We selected 3D object detection as a representative example of downstream tasks. To evaluate robustness across different perception architectures, we consider three representative LiDAR detectors: PointPillars (BEV-based), SECOND (voxel-based), and PointRCNN (point-based).
Fig. 5 shows the Car 3D bounding-box AP@0.7, 0.7, 0.7 as a function of bitrate for the original and reconstructed point clouds, evaluated with PointPillars, SECOND, and PointRCNN. The dashed line indicates the detection accuracy on the uncompressed point clouds. The results demonstrate that the proposed method achieves higher detection accuracy than the baselines across all three detectors in low-bpp regimes, i.e., bpp from 1.0 to 2.0.
@@ IV-B3 DECODING LATENCY
Table II shows the average decoding latency of the proposed and baseline methods for LiDAR frame 00 of sequence 00. The decoding latency values for the proposed method and the baselines are the total time required from RI decoding to the 2D-to-3D mapping. The decoding latency of the proposed method is comparable to that of image compression methods and has more than 65.5% and 93.3% reduction compared to G-PCC, Draco and R-PCC methods, respectively. In addition, the proposed scheme achieves a speedup of over four orders of magnitude compared to OctAttention. This means that the proposed method approaches the decoding latency of image compression methods and achieves 3D reconstruction quality comparable to PCC/R-PCC methods.
IV-B4 ENCODING LATENCY
Table III lists the average encoding latency of the proposed and baseline methods for LiDAR frame 00 of sequence 00. Here, the encoding latency for the RI-based schemes contains the conversion time from the point cloud to the RI. It can be seen that INR-based approaches, including the proposed one, involve significantly longer encoding time than both 3D point cloud and 2D image compression methods. However, as shown in Table II, INR-based compression drastically reduces decoding latency, which is advantageous for on-demand, quality-driven services.
We note that the proposed scheme has a trade-off between reconstruction quality and encoding latency depending on the learning rate schedule. Here, we use initial and minimum learning rates of $1\times 10^{-4}$ and $1\times 10^{-12}$ for quality-sensitive users. When we use initial and minimum learning rates of $1\times 10^{-6}$ and $1\times 10^{-8}$ for encoding, the encoding latency is reduced to 30 minutes with a slight degradation in reconstruction quality. This means the proposed scheme can select quality-oriented or latency-oriented configurations based on application requirements.
((a))
((b))
@@ IV-B4 ENCODING LATENCY
Refer to caption ((a)) Refer to caption ((b))
Refer to caption ((c)) Refer to caption ((d))
Figure 7: Snapshot of the reconstructed LiDAR point clouds in pixel-wise and patch-wise proposed methods. Here, (a)-(b) and (c)-(d) show the reconstructed point clouds of frames 00 and 25 of sequence 00, respectively.